# StyleMaster: Towards Flexible Stylized Image Generation with Diffusion Models

### v1

@@ 6 Conclusion
We try to solve Stylized Text-to-Image Generation in this paper. A novel model is proposed to solve the problems of insufficient style and inconsistent semantics which are suffered by previous methods such as StyleAdapter. The improvement mainly comes from the multi-source style embedder, in which multiple sources and used to achieve comprehensive style embeddings and eliminate the semantic information from style reference images, and the dynamic attention adapter, in which style embeddings dynamically interact with attention layers in diffusion UNet. Extensive experiments have been conducted to show the efficacy of our proposed method as a powerful stylization method, which can be widely applied to real-life scenarios.
Appendix A Experiment details
@@
In order to make some results (Fig. 1 and Fig. 6) in main paper simple and easier to understand, we omit the style reference images and summarize the used styles here:
  • Fig. 1 from first row to fifth row: pixar, ink, wooden, cubism, Van Goah.
  • Fig. 6 first row from left to right: Degas, Cezanne, Cubism.
  • Fig. 6 second row from left to right: ink, flat cartoon, pencil.
  • Fig. 6 third row from left to right: flat cartoon, impasto, watercolor.
@@
Prompts used in figures in main paper.
All qualitative results in the qualitative ablation study of main paper use the same prompts as in the quantitative evaluation, concretely as follows,
  • Fig. 6 first row from left to right: A girl wearing a red dress, she is dancing; A little boy with glasses and a watch; A smiling little girl.
  • Fig. 6 second row from left to right: A puppy sitting on a sofa; An curly-haired boy; A boy wearing glasses, he is reading a thick book.
  • Fig. 6 third row: A man wearing a black leather jacket and a red tie.
Appendix B More discussion
@@
We would like to highlight two limitations of our method. First, the multi-source style embedding extraction model relies on a patch-level transformer. In this way, it is less efficient for this model to process too many style reference images. While such a scenario is to some extent unrealistic, since it is generally sufficient to represent a specific style with less than 10 images, solving this problem can be related to improving vision transformer structures, which can be taken as future works. Second, the proposed method is only available for style conditions in the form of images. Other forms such as texts, videos and 3D data are not considered in this work and can be solved in the future.
@@
To show that our method is generalizable enough, we further apply our method to pretrained SDXL, which is an advanced version of SD. The results are shown in Fig. 9 and Fig. 10. We can find that basically the proposed method can introduce correct style to pretrained SDXL. Since there is significant gap between the prior knowledge learned by SDXL and SD1.5, the generated images also show different patterns. The results show that SDXL can better handle styles regarding lines and colors, while SD1.5 can provide better global-level styles. Moreover, SDXL can make better balance between style and general image aesthetic. The human faces generated by SDXL are more proper, thanks to its larger capacity.
Figure 9: Quantitative results of our proposed method with SDXL in one-shot setting. Styles used in each row from up to bottom: Cezanne, flat cartoon, cubism, pencil.
@@
Effectiveness of frequency domain decomposition.
In the multi-source style embedding extraction module, we utilize the discrete wavelet transform to first decompose the patch level features of style reference images into low-frequency and high-frequency features. To see how this process can help our module, we visualize the attention weights for two sets of reference images among different denoising time steps in Fig. 11. Three main phenomenon can be concluded: (1) Style embeddings concentrate more on low frequency features, which is reasonable since low-frequency features contain information such as color. (2) The patterns of feature usage are consistent among two prompts for each style, while being different for different styles. (3) For high-frequency features, different timesteps generally focus on different information. These results can sufficiently support our design of decomposing the image features regarding frequency.
Quantitative results for ablation study.
We present the quantitative results for all variants used in Tab. 5, which are generally consistent with the qualitative results in the main paper. Interestingly, we find that when not using dynamic adaptor, the model has a very different tendency compared with the full model, with much better text prompt similarity but much worse style similarity. This is reasonable since the proposed dynamic adaptor greatly strengthens the impact of style embedding during both self-attention and cross-attention. In this way, the cross-attention with text prompts is weakened is disguise. In general, adopting dynamic adaptor can make a good balance between text prompt fidelity and style fidelity.
Table 5: Quantitative ablation study for multi-shot setting. For all metrics, the larger score denotes the better model.
multi-shot Methods Text Sim StyleSim
w/o VGGToken 0.288 0.698
w/o NegEmb 0.281 0.696
w/o OnlyUP 0.286 0.702
w/o dynamic adaptor 0.299 0.643
w/o style loss 0.290 0.705
w/o disen loss 0.286 0.694
Ours 0.291 0.719
Image-to-image.
Apart from the basic text-to-image task, we also extend our method to image-to-image task, in which we follow the commonly used pipeline to first and then adopt the ControlNet-Canny [24] together with our proposed method to generate a new image with target style. The results are shown in Fig. 12. One can find that when given different reference images such as Japan anime, American anime, Von Goah’s painting and pixar anime, the human face in the base image can flexibly change according to the style, which illustrates the effectiveness of our method.
Figure 12: Image to image results generated by our model.
Appendix C More quantitative results
@@
Figure 15: More quantitative results of our proposed method in multi-shot setting. Styles used in each three rows from up to bottom: surreal, psychedelic, surf.
We provided more quantitative results in both one-shot and multi-shot settings in Fig. 13, Fig. 14 and Fig. 15. The images generated with each group of reference images share consistent style, while correctly showing the target objects.

### v2

@@ 6 Conclusion
We try to solve Stylized Text-to-Image Generation in this paper. A novel model is proposed to solve the problems of misinterpreted style and inconsistent semantics which are suffered by previous methods such as StyleAdapter. The improvement mainly comes from the Mixed Style Descriptor, in which multiple sources and used to achieve comprehensive style embeddings and eliminate the semantic information from style reference images, and the dynamic attention adapter, in which style embeddings dynamically interact with attention layers in diffusion UNet. Extensive experiments have been conducted to show the efficacy of our proposed method as a powerful stylization method, which can be widely applied to real-life scenarios.
7 Experiment details
@@
In order to make some results (Fig.1 and Fig.6) in main paper simple and easier to understand, we omit the detailed style reference images and summarize the used styles here:
  • Fig.1 from up to bottom: america cartoon, wooden, cubism, Van Goah.
  • Fig.6 from up to bottom: pencil, watercolor, monet, impasto.
  • Fig.7 from up to bottom: Degas, ink.
  • Fig.8 from up to bottom: Cezanne, impasto.
  • Fig.9 from up to bottom: pencil, watercolor.
@@
8 More discussion
@@
We would like to highlight two limitations of our method. First, the mixed style descriptor (MSD) relies on a patch-level transformer. In this way, it is less efficient for this model to process too many style reference images. While such a scenario is to some extent unrealistic, since it is generally sufficient to represent a specific style with less than 10 images, solving this problem can be related to improving vision transformer structures, which can be taken as future works. Second, the proposed method is only available for style conditions in the form of images. Other forms such as texts, videos and 3D data are not considered in this work and can be solved in the future.
@@
Comparison with InstantStyle-SDXL and StyleAlign.
In Fig. 12 we present the comparison between our method and InstantStyle and StyleAlign, both using SDXL as backbone network. We find that while InstantStyle does perform better with SDXL than SD1.5, suffering less from content leakage, it tends to generate images with classic art styles such as paintings. This makes the generated results less similar to the reference images. On the other hand, InstantStyle-SDXL still cannot handle styles such as cubism and impasto, which can be well modelled by our method. Moreover, StyleAlign is generally worse than the other two methods.
Figure 12: Qualitative comparison with InstantStyle and StyleAlign using SDXL as backbone.
Effectiveness of frequency domain decomposition.
In the mixed style descriptor (MSD), we utilize the discrete wavelet transform to first decompose the patch level features of style reference images into low-frequency and high-frequency features. To see how this process can help our module, we visualize the attention weights for two sets of reference images among different denoising time steps in Fig. 13. Three main phenomenon can be concluded: (1) Style embeddings concentrate more on low frequency features, which is reasonable since low-frequency features contain information such as color. (2) The patterns of feature usage are consistent among two prompts for each style, while being different for different styles. (3) For high-frequency features, different timesteps generally focus on different information. These results can sufficiently support our design of decomposing the image features regarding frequency.
@@
To show that our method is generalizable enough, we further apply our method to pretrained SDXL, which is an advanced version of SD. The results are shown in Fig. 17. We can find that basically the proposed method can introduce correct style to pretrained SDXL. Since there is significant gap between the prior knowledge learned by SDXL and SD1.5, the generated images also show different patterns. The results show that SDXL can better handle styles regarding lines and colors, while SD1.5 can provide better global-level styles. Moreover, SDXL can make better balance between style and general image aesthetic. The human faces generated by SDXL are more proper, thanks to its larger capacity.
Figure 14: Multi-shot qualitative results with non-related objects added to the caption of style reference images. Prompts for every two columns from left to right: A bird in a word; A daisy with a ladybug on it; A stone with a crack in it, holding a plant growing out of it; A puppy sitting on a sofa.
Reasonableness of negative semantic embedding.
One would ask whether it is proper to directly subtract style token part in $\hat{z}_{caption}$ from $\hat{z}_{CLIP}$ and if the subtracted vector could inadvertently contain elements of the negative prompt (e.g., “reading a book”), rather than purely style information. Note that after each subtraction an attention is further applied to $\hat{z}^{CLIP}$, where unrelated negative prompts would be weakened. To verify this we provide several examples on multi-shot ink and watercolor styles. Specifically, we add a non-related caption ‘There is a robot, a UFO and a monster in the image.’ to the caption of each reference image. The results are presented in Fig. 14, in which the semantics of generated images do not degrade compared with original ArtWeaver.
Figure 15: Qualitative results of ArtWeaver under 10-shot setting. The first and third rows contain the reference images, the other rows show the generated results.
Image-to-image.
Apart from the basic text-to-image task, we also extend our method to image-to-image task, in which we follow the commonly used pipeline to first and then adopt the ControlNet-Canny [28] together with our proposed method to generate a new image with target style. The results are shown in Fig. 16. One can find that when given different reference images such as Japan anime, American anime, Von Goah’s painting and pixar anime, etc. The human face in the base image can flexibly change according to the style, which illustrates the effectiveness of our method.
Figure 16: Image to image results generated by our model.
Results with more shots.
To further show the versatility of our proposed method, we conduct a 10-shot experiment, of which the results are presented in Fig. 15. The results show that our method, when given reasonable reference images, is robust and well performing under 10-shot setting.
@@
9 More qualitative results
@@
We provided more quantitative results in both one-shot and multi-shot settings in Fig. 18, Fig. 19. The images generated with each group of reference images share consistent style, while correctly showing the target objects.